OpenStudy (anonymous):

simplify 1/x-1 I think the answer is 1 but i need explanation why please

OpenStudy (anonymous):

so is it (1/x) -1 or 1/(x-1)

OpenStudy (anonymous):

oh wait its actually 1/(x^-1) my bad.

OpenStudy (anonymous):

oh ok. so why do you think it is 1?

OpenStudy (anonymous):

Well theres a rule for it i think for example 6^3 in a rational it'll be 1/(6^3) which itll eventually be 1/216 i think that applies to this problem too right?

OpenStudy (anonymous):

Well, lets think about a simpler example. Do you know what 2^-1 would be?

OpenStudy (anonymous):

1/2 right?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so what about 1 / 2^-1

OpenStudy (anonymous):

1 / -2 but since its in the denominator its already considered a negative number i believe.

OpenStudy (anonymous):

its not negative 1/2

OpenStudy (anonymous):

There is no exponent that will make a number become negative

OpenStudy (anonymous):

and a number being in the denominator doesn't mean its negative either

OpenStudy (anonymous):

Ohh alright i understand that, although this specific property confuses me a lot.

OpenStudy (anonymous):

yeah it can do that

OpenStudy (anonymous):

so 1 / 2^-1 would be 1 / 1 / 2

OpenStudy (anonymous):

and one divided by one half is two

OpenStudy (anonymous):

So it's 2?

OpenStudy (anonymous):

so when you have a negative exponent in the denominator, you are basically just moving it back to the numerator.

OpenStudy (anonymous):

just like when you have a negative exponent in the numerator, you move it to the denominator

OpenStudy (anonymous):

yes its 2

OpenStudy (anonymous):

Hmm but the question then asks my why x^1/3 * x^1/3 * x^1/3 is the same as that question. The answer for that one would be x^3/3 which would equal 1 right?

OpenStudy (anonymous):

(x ^ (1/3)) * (x ^ (1/3)) * (x ^ (1/3)) = x

OpenStudy (anonymous):

when you multiply some base number exponents, so long as the base in the same you add the exponents

OpenStudy (anonymous):

so you're just adding (1/3)+(1/3)+(1/3) = 1. But x^1 = x, not 1.

OpenStudy (anonymous):

Well yeah i meant that but then how is that the same as the previous one?

OpenStudy (anonymous):

Because the answer is the same.

OpenStudy (anonymous):

They're both equal x

OpenStudy (anonymous):

1/x^-1 = x

OpenStudy (anonymous):

Ohhhh they're equal to x?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ahh may i ask you the other two then?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Its an add on to the question. This one is 11√x^5*x^4*x^2, i put that it turns out to be 11√x^11 which will then make it 11/11=x which is 1=x

OpenStudy (anonymous):

right. The 11th root will undo the 11th exponent leaving only x.

OpenStudy (anonymous):

So I'm right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome and now the last one is 3√x^3 which i put that it'll make 3/3 which is also equal to x.

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

Awesome. There's a really confusing one here that has to do with the previous one. If you don't mind that'll be my last one?

OpenStudy (anonymous):

Ok lets see

OpenStudy (anonymous):

Simplify 1/(3√x-6) I put that it would make 1/(-6/3) which it will make it be 1/-2 = -1 i stopped there.

OpenStudy (anonymous):

Ok, so what you want to do here is multiply both the top and bottom by something that will get rid of the radical in the denominator

OpenStudy (anonymous):

we have the 3rd root of x-6

OpenStudy (anonymous):

or (x-6)^(1/3)

OpenStudy (anonymous):

oh man sorry its actually x^-6

OpenStudy (anonymous):

I always forget the exponent sign. sorry

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

So 1/ \[\sqrt[3]{x^-6}\]

OpenStudy (anonymous):

sorry, thats messed up looking, but is that right?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

ok. So you can rewrite this as (\[\sqrt[3]{x^6}\]

OpenStudy (anonymous):

Without the -6 exponent?

OpenStudy (anonymous):

or x^(6/3)

OpenStudy (anonymous):

right, because we bring it into the numerator

OpenStudy (anonymous):

It makes more sense when you look at everything with fractional exponents rather than radicals

OpenStudy (anonymous):

Ohhhhh wow that makes sense! I see now and so if it where x^6/3 wont it be x^2?

OpenStudy (anonymous):

because you had 1 / x^(-6/3)

OpenStudy (anonymous):

originally

OpenStudy (anonymous):

Yeah i see it's a lot easier to see it that way in a fraction.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So is it x^2?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thanks man i appreciate it very much

OpenStudy (anonymous):

np. thanks for actully trying to learn lol

OpenStudy (anonymous):

Haha no problem, better now than never.

OpenStudy (anonymous):

wait until you get to logarithms if you think dealing with exponents llke this is confusing

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks a lot. good luck